### v1

# Travel Experience in Public Transport Dataset

Abstract
The transportation sector holds the potential to change the world towards a greener future if it aligns with increasing mobility needs. One solution is to make public transport an attractive alternative to individual transportation. Real-world data is needed to investigate reasons for and indicators of positive and negative travel experience. Here, we present a GPS-tagged dataset where participants wore an electrocardiogram and reported experience sampling that measured stress, satisfaction, events, and emotions while traveling by tram, train, and bus. An interactive experience map helps to visually explore the data. As benchmark analysis for future users of the dataset, we report significant stress hot spots and satisfaction cold spots during the participants’ journeys. The reported events and emotions, especially in such hot and cold spots, can be analyzed to highlight points of positive and negative travel experience in an ecologically highly valid setting. Data on age and self-identified gender offers insights to differences between user groups. Overall, our dataset enables the combination of qualitative and quantitative methods to identify user’s needs in public transportation.
Background & Summary
Transport causes nearly 25% of the world’s carbon dioxide emissions from energy [1]. Buses, railways, and trams constitute 13.8% of kilometers driven by motorized transport on land in the EU in 2020 [2]. The other 86.2% of kilometers were covered by cars [2]. At the same time, public transport transmits more than 50% less greenhouse gases per passenger kilometer than cars when compared to small and medium cars only [3]. One effective step towards worldwide greenhouse gas reduction is, therefore, a shift from private to public transport [4].
While the importance of shifting to public transport is evident, the feasibility of this shift has been the focus of extensive studies. Therefore, a large body of research has investigated factors to enable an increase in public transport usage in terms of efficiency [5, 6], infrastructure [7], and politics [8]. Another important factor for a transition towards greener mobility is the user, who is the ultimate deciding factor in increased public transport usage. Therefore, previous studies have mostly concentrated on time, cost, and reliability aspects that relevantly influence mobility choices ([9, 10, 11]).
Despite the extensive research on technical and practical aspects, the human element - specifically the travel experience and its influence on future travel mode choices - has not been thoroughly explored. Travel experience is closely linked with feelings and emotions [9]. Such feelings and emotions are experienced through a person’s cognitive appraisal and an appropriate physiological response to an event [12]. It is, therefore, essential to use multimodal metrics that consider both the traveler’s emotional state and physiological response [13]. Thus, when assessing travel experiences in intermodal journeys, it is beneficial to integrate subjective assessments with physiological indicators of emotions, as well as relevant contextual data [14, 15]. Other indicators of emotion, like facial expressions or electroencephalography, are not feasible in public transport. While numerous studies have used physiological data to correlate physiological changes with subjective experiences and emotions [16, 17, 18], only first steps have been taken to use physiological data to infer travel experience in public transport [19, 20].
Experience sampling is a method to collect subjective data with high frequency in a ecologically valid setting [21]. By capturing moments of positive and negative experiences as they occur, researchers are provided with data that is less influenced by memory biases in terms of quantity and intensity compared to post-hoc methods [22]. The method has previously been used to assess information need in an automated shuttle [23] and, as a case study, in public transportation of Santiago de Chile [14]. A straightforward method to depict such location-based experience sampling data is through map visualizations. These have previously been used in the automotive sector [24, 25] and for pedestrians and cyclists [26]. Using experience mapping to investigate public transport experience has been suggested by [27] as a method that can pinpoint collective points of negative experience in a transportation system based on the first 23 participants of this dataset. In this paper, we present the complete dataset and include a map visualization.
Overall, the qualitative dimension of the travel experience has been addressed in some studies [28, 29], but synchronous, data-driven datasets are lacking so far. Given this gap, our dataset offers a comprehensive and synchronous approach by collecting data from 44 participants who traveled on a pre-defined route that entailed changes between tram, bus, and train on a 15 km long route. During the journey, we collected global positioning system (GPS) data, heart rate and heart rate variability data, and experience sampling data on satisfaction with travel, stress levels, freely chosen emotion words, and events that occurred. We visualized all these data in an interactive map that enables the researcher to easily investigate the dataset. We identified significant spots of stress and satisfaction and the events and emotions that were reported at these spots as benchmark analysis.
The dataset can be used to develop data analyses methods related to causes for positive and negative travel experience, locations of collective positive or negative travel experiences, and physiological correlates of travel experience during active and passive travel modes. Data on age and self-identified gender can be used to investigate differences between user groups. Overall, our dataset enables the combination of qualitative and quantitative methods to identify methods that measure user’s needs in public transportation.
Methods
Procedure
Participants arrived at our institute and were informed about our study’s procedure (see Figure 1) and aim. They were equipped with a cell phone and an electrocardiogram (ECG) belt, which they could keep for up to three weeks. Within this time, all participants traveled the same predefined route in both directions between three to five times, depending on the participant’s availability. They recorded their way from and to the research institute, where they were free to choose a transportation mode. The one-way route consisted of
  • Getting from home with a freely chosen transportation mode to a central bus station in Brunswick (German city with roughly 250,000 inhabitants),
  • taking a tram to Brunswick main station (GPS location: 52.251332328, 10.537164518),
  • taking a bus (50% of journeys) or a train (the other 50% of journeys) to Wolfenbuettel (German city with roughly 50,000 inhabitants, distance roughly 15km, GPS location: 52.15951774394025, 10.532477626246047),
  • walking to the city center.
and was taken back in reverse. Participants were free to choose to start their journey from Brunswick or Wolfenbuettel. The route was chosen to include as many as possible transportation modes of public transport. Figure 2 displays the route. Data acquisition took place from January to December 2023.
Participants
In total, 44 participants (22 self-identified as male, 20 self-identified as female, and two preferred not to fill in any demographic information) took part in the study. They had a mean age of 29.4 years (SD 13.8 years), and none of them indicated to have had any cardiovascular diseases. They indicated to have no specific restrictions during travel except for two persons who indicated to travel with a lot of luggage on all of their journeys. One participant lived in a city with over 500,000 inhabitants, 34 participants lived in a city with up to 500,000 inhabitants, two up to 100,000, four up to 50,000, and one in the countryside. Eighteen had a way to work below 5 km, six had 10-20km, three had 20-30 km, two 30-40km, nine 5-10km, and four less than 1 km. Thirty-four of the participants had a driver’s license. With multiple responses possible, 28 participants usually take a bus to work, 8 drive to work by car, 1 is taken to work by car, 23 cycle to work, 12 take a train to work, 16 a tram and 19 usually walk to work. The participants were acquired from our institute’s participant pool and received 60€-100€ (depending on the number of journeys taken, 10€ per journey) as financial reimbursement for participation. Before starting the study, participants were informed about and agreed to the study’s procedure and the handling of personal data (in line with the General Data Protection Regulation of the EU). The study’s procedure and sharing the data was approved by the institution’s ethical committee (reference no. 07/22).
Measurements
Subjective measures:
Before the start of the study, participants filled in a demographic questionnaire and the Big Five Inventory (BFI-10) following [30] on a scale from 1 = strongly disagree to 5 = strongly agree with the following questions:
  • "I am rather reserved" - scores on Extraversion, reversed
  • "I am generally trusting" - scores on Agreeableless
  • "I tend to be lazy" - scores on Conscientiousness, reversed
  • "I am relaxed and handle stress well" - scores on Neuroticism, reversed
  • "I have few artistic interests" - scores on Openness, reversed
  • "I am outgoing, sociable" - scores on Extraversion
  • "I tend to find fault with others" - scores on Agreeableless, reversed
  • "I do a thorough job" - scores on Conscientiousness
  • "I get nervous easily" - scores on Neuroticism
  • "I have an active imagination" - scores on Openness
They also filled in the Affinity to Technology (ATI-9) scale following [31] that is computed as mean over all 9 scores on a scale from 1 = ’completely disagree’ to 6 = ’completely agree’ with the following questions:
  • "I like to occupy myself in greater detail with technical systems"
  • "I like testing the functions of new technical systems"
  • "I predominantly deal with technical systems because I have to" - reversed
  • "When I have a new technical system in front of me, I try it out intensively"
  • "I enjoy spending time becoming acquainted with a new technical system"
  • "It is enough for me that a technical system works; I don’t care how or why" - reversed
  • "I try to understand how a technical system exactly works"
  • "It is enough for me to know the basic functions of a technical system" - reversed
  • "I try to make full use of the capabilities of a technical system"
At the beginning of every journey, participants indicated their destination and were asked to indicate every mode change of their journey by pressing the corresponding button in the MovisensXS App (Movisens GmbH in Karlsruhe, Germany, 2023b) on the study smartphone. During their journey, participants were prompted to answer an experience sampling questionnaire every 3.5 minutes using the same app. The decision to adopt a 3.5-minute interval was based on a previous study [20, 32], during which participants were exposed to different frequencies of questionnaire requests. Participants used Likert scales ranging from 1 (indicating low stress and low satisfaction) to 10 (reflecting high stress and high satisfaction) to assess their current satisfaction with the journey and their stress levels. The precise questions asked as translated from German were ’Please rate your current satisfaction with travel’ and ’how stressed do you feel?’. They furthermore answered ’how do you feel independent of your journey?’ from 1 (positive) to 10 (negative), ’How would you describe your current emotion?’, and ’Did anything happen since the last journey?’. If answered with ’yes’, they could choose from the following event categories that were found in a prestudy [20, 32]:
  • "Delay / Connection missed / Hurry required," (will be referred to as ’Delay’)
  • "Disturbing people in the environment / crowded transportation," (will be referred to as ’Disturbing People’)
  • "Driving behavior of the driver," (will be referred to as ’Driving Behavior’)
  • "Infrastructure (e.g. no roofing/no or not enough information /not barrier-free / uneven road)," (will be referred to as ’Infrastructure’)
  • "Positive social situation/interaction," (will be referred to as ’Positive Social Interaction’)
  • "Media entertainment (music, podcast, book, etc.)," (will be referred to as ’Media Entertainment’)
  • "Destination or connection reached / means of transport on time," (will be referred to as ’Destination/connection reached’)
  • "Discomfort (boredom, impatience, difficult thoughts, hunger, fatigue, cold, wetness, etc.).," (will be referred to as ’Discomfort’)
  • "Transportation is comfortable (cozy/pleasant temperature / not too crowded / get a seat, etc.)," (will be referred to as ’Comfort’)
  • "Beautiful surroundings,"
  • and an option to "report a different category." (will be referred to as ’Others’)
The participants were furthermore free to report a freely chosen emotion word below the choice of categories. We translated the free text inputs to English with Deepl.com. Between the prompted questionnaires, the participants were instructed to freely prompt another questionnaire whenever they encountered something that changed their evaluation or experience of the commute. At the end of every journey, participants rated their overall travel experience using the Satisfaction with Travel Scale (STS) [33]. The STS assesses the emotional and cognitive evaluation of the journey. It comprises of the questions ’How do you feel?’ from Likert scales that measured hurried (1) to relaxed (7), worried (1) to confident (7), stressed (1) to calm (7), tired (1) to alert (7), bored (1) to enthusiastic (7) and annoyed (1) to stimulated (7). Part two of the STS asks ’How do you judge your journey?’ from ’the worst journey I can imagine’ (1) to ’the best I can imagine’ (7), from ’low quality’ (1) to ’high quality (7), and from ’the journey worked out badly’ (1) to ’the trip went smoothly’ (7).
After all journeys, the participants filled out the post-questionnaire containing two decision questions:
  • Will you continue to commute using public transport?
  • If you had to travel from Wolfenbuettel main station to Brunswick main station (or vice versa), would you prefer using the bus or the train?
Heart Rate Measures:
The participants’ heart rate and heart rate variability (HRV) was once per second extracted from an EcgMove 4 belt from Movisens (Movisens GmbH, Karlsruhe, Germany, 2022).
Context Data:
Throughout their travel, the location was consistently recorded using the movisensXS App and the Movinglab app  [^1] https://movinglab.dlr.de/. While the movisens XS app collected subjective and GPS data, the Movinglab app automatically recognized the travel mode (walk, stationary, train, change, bus, car, bicycle, unknown) and also recorded GPS data. To adhere to the guidelines set by our institute’s data protection committee, participants received explicit instructions to activate the GPS data on the study cell phone solely upon entering the initial major street subsequent to commencing their journey. The GPS signal had a mean accuracy of 8.60m.
Data Preprocessing
The original heart rate data (column HR [1/min]_EcgMove4) was transformed using the median absolute deviation per participant (column ’HR_mad)’ by 1) centering the values by subtracting the participant’s median heart rate from all heart rate data, 2) calculating these values’ median absolute deviation and 3) division of the centered values by the median absolute deviation values (see 0_mad_per_participant.Rmd of the available code). Furthermore, we excluded heart rate values during times that the ECG belt recognized active travel modes like walking or standing, and only included data where participants were sitting for longer than one minute in the column HR_mad_filtered.
Data Records
The data is available on the Open Science Framework [34]. Each .csv contains the data of one participant. Find each column explained in the tables at the end of his paper.
Technical Validation
We include an interactive map  [^2] https://estherbosch.shinyapps.io/affectivemaps that enables dynamic visualization of our dataset with the R Shiny package [35] in R [36]. The data displayed on the map were filtered to only include data where more than two participants contributed to an area. This was done by rasterizing the data into 100mx100m longitude and latitude squares and excluding all squares that contained less than two participants, (see 1_HR_mad_filtered_2squares.Rmd). All relevant variables measured during the journeys can be chosen by a drop-down menu and data of all participants or single participants can be selected, see Figure 3. A click on each data point provides all available information on this point’s stress rating, satisfaction rating, heart rate, and freely reported emotions.
Cold-/Hot Spots
As reference analysis for detecting relevant points of travel experience in public transport with our dataset, we calculated cold spots and hot spots of satisfaction and stress ratings using the Getis-Ord Gi statistic following [37] with 100m radius circles around each data point. This is an analysis to indicate a statistically significant cluster of high or low values in each rating. A statistically significant hot spot exhibits a high value and has neighboring high values. The cumulative value of a feature (the data points within the circle around a data point with a 100m radius, in this case) and its adjacent data points are assessed relative to the cumulative value of all features. When this cumulative value significantly diverges from the anticipated local sum more than expected by random chance, it represents a statistically significant stress or satisfaction rating. To all calculations, we applied Bonferroni correction for multiple tests. The hot and cold spots are displayed on the interactive map as the middle points of the 100m radius circles with statistically significant hot and cold spots. The respective hot- and cold spot reported events within the 100m radius are displayed in Figure 4. There were 28 freely reported emotions in the satisfaction cold spot, of which 11 (39.3%) were positive ("calm", "happy", "joy", "relaxed", "satisfied", "in a good mood and relaxed", "motivated", "excited", "good", "serene", "enjoyable"), also 11 (39.3%) were negative ("upset", "annoyed", "tired and funny to drive past home", "tired", "slightly tense", "tense", "a little annoyed and tired", "annoyed but getting better", "still annoyed", "slightly annoyed", "bored") and 6 (21.4%) neutral ("okay", "neutral", "indifferent", "unchanged", "slow", "quiet"). There were 12 freely reported emotions in the stress hot spot, of which 3 (25%) were positive ("Good","Relaxed", "Satisfied"), 4 (33.33%) were negative ("Stressed", "Tired", "Stressed because the tracked mode does not show up","Annoyed") and 5 (41.67%) neutral ("Quiet", "Okay", "Good, a little stressed because late, a lot to do", "Talkative", "Neutral").
Heart Rate Measures and Permutation Analysis of Heart Rates in Cold-/Hot Spots
[…]
Usage Notes
We recommend to use the interactive map to get an overview of the data. We also recommend to use Heart Rate and HRV only filtered by physical activity through transportation mode, with one suggestion being the filtered heart rate we used in the exemplary analyses. To more reliably know whether a participant was physically active we recommend the GPS data, the self-reported mode and the transportation mode detected by the Movinglab app. Furthermore, note that GPS data has a mean accuracy of 8.60m. Not all Experience Sampling data (variables ending with _xs) have a corresponding Longitude and Latitude value, so that filtering by valid Longitude and Latitude values excludes some of the freely reported experience sampling values. This is, therefore, the case for all dataframes but the original ’ExperienceMapsDataset.csv’ when running the code. Both the subjective travel mode report (ModeButton) and the automated travel mode recognition (Mode_keepmovingLast) should be treated with caution, as the respective GPS location indicates that both are not reliable (i.e., participants forgot to indicate their mode change and the Movinglab app sometimes recognized wrong travel modes). The data have limited location variation with the included German cities, and the results’ generalizability needs to be evaluated in future studies.
Code availability
All code (written in R) used to process our data as explained in this paper (Preprocessing, Experience Maps, Hot and Coldspots Analysis) can be found in the same folder as the dataset.